# Conformation and dynamics of wet tangentially-driven active filaments

### v2

@@ II.5 Simulation parameters
We choose the monomer diameter $\sigma$ as the unit of length, $k_{B}T$ as the unit of energy and $\tau=\sqrt{m\sigma^{2}/k_{B}T}$ as the unit of time. The wet polymers are simulated in a three-dimensional box with sides of length $L_{box}=60$ and periodic boundary conditions. The simulation box is split into collision cells of size $L_{cell}=1$. The mass of the MPC fluid particles is set to $m=1$ and their collision time step is set to $\delta t=0.01$. The stochastic rotation angle is chosen as $\alpha=130^{\circ}$. The number density of the MPC particles is $\langle N_{c}\rangle=10$ leading to a total of $2.16\times 10^{6}$ MPC particles in the simulation box. The viscosity of the solvent made of MPC particles is given by $\eta=82.1\sqrt{\frac{mk_{B}T}{L_{cell}^{4}}}$ Winkler and Huang (2009). Therefore, the chosen parameters create a viscosity of $\eta=82.1$.
@@
We simulate 150 individual wet polymers. For each polymer, we first generate a three dimensional random walk of length $N_{m}$ and step size $\ell$. We use this sequence as the initial positions of the monomers. Once the polymer is created, the system is simulated for $10^{7}$-$10^{8}$ time steps to reach a steady state depending on the value of $f^{a}$. Then we run a production simulation of $2\times 10^{7}$ $2\times 10^{8}$ time steps, depending on the activity level. Simulations of dry active polymers, conducted for comparison with their wet counterparts, involve 120 individual dry active polymers described by Eq. (15) in a box with size $L_{box}=200$, while all other parameters being identical to those of wet active polymer model. The friction coefficient of monomers in dry active polymer simulation is chosen to be $\gamma=3\pi\eta\sigma\approx 773$. This leads to a ratio of $M/\gamma\approx 0.013$ which ensures that the dynamics of monomers are overdamped and inertial effects Fazelzadeh et al. (2023) remain negligible. To reach the steady state, we use $10^{7}$ to $10^{9}$ time steps, depending on the activity level. For the production run, twice the number of time steps used to reach the steady state are used. The time step for dry simulations is $\Delta t^{\prime}=0.002$.
@@ III Conformational features of wet active polymers
Figure 4: The probability distribution function of the end-to-end distance $P(R_{e})$ for different active forces $f^{a}$ for chains with $\kappa=0$ and $10$. Solid and dashed lines represent data for chains with and without hydrodynamic interaction, respectively.
@@ III.1 End-to-end distance
To have a quantitative assessment of hydrodynamic effects on the mean size of active polymers, we examine the root of the mean squared end-to-end distances of the chains, i.e., $\sqrt{\langle R_{e}^{2}\rangle}=\sqrt{\langle|\mathbf{\bm{R}}_{N_{m}}-\mathbf{\bm{R}}_{1}|^{2}\rangle}$. Fig. 3 depicts the $\sqrt{\langle R_{e}^{2}\rangle}$ against activity for dry and wet chains with $\kappa=0$ and $10$. For flexible chains with $\kappa=0$, we observe slightly smaller mean end-to-end distance values for wet polymers compared to their dry counterparts. As the activity level increases, the difference between the two becomes more pronounced. Looking into the $\sqrt{\langle R_{e}^{2}\rangle}$ of semiflexible chains with $\kappa=10$ in Fig. 3, we observe a more notable shrinkage of chains compared to their flexible counterparts, specially at larger active forces $f^{a}>1$ where the mean size of wet active polymers is about twice smaller than the dry ones.
To better assess the impact of HI on the conformation of active polymers, we examine the probability distribution function (PDF) of $R_{e}$, presented in Fig. 4. For both flexible and semiflexible polymers, the distribution of end-to-end distance of wet active polymers shifts towards smaller $R_{e}$ values relative to their dry counterparts and the $P(R_{e})$ of wet active chains become narrower. This effect becomes more pronounced with an increase in active force, particularly noticeable in the probability distributions $P(R_{e})$ of wet active polymers at $f^{a}=25$. In both flexible and semiflexible chains, narrow peaks emerge at lower $R_{e}$ values for wet active polymers, contrasting sharply with their dry counterparts. In the case of semiflexible chains, the HI-induced shrinkage of conformation is substantial enough to alter the overall shape of wet active polymers from a polymer-like conformation similar to that of dry polymers into a helix-like conformation, as demonstrated in Fig. 2.
Refer to caption(a)(b)(c)(d)
Figure 5: The mean squared internal distance (MSID) $\langle R^{2}(s)\rangle$ versus curvilinear distance $s$ for wet and dry polymers with bending stiffness of $\kappa=0$ and active forces of a) $f^{a}=0.1$ and b) $f^{a}=25$ and bending stiffness of $\kappa=10$ with active forces of c) $f^{a}=0.1$ and d) $f^{a}=25$.
@@ III.2 Interachain correlations
Having investigated the effects of HI on the statistical properties of end-to-end distance of active polymers, we study its impact on intrachain correlation functions by examining the mean squared internal distance (MSID) and bond-bond correlation function.
First, we examine the MSID of the chains defined as $\langle R^{2}(s)\rangle=\langle(\mathbf{\bm{R}}_{i}-\mathbf{\bm{R}}_{i+s})^{2}\rangle$ presented in Fig. 5. At the low activity level of $f^{a}=0.1$, MSIDs of wet chains are slightly smaller than their dry counterparts presented Fig. 5 (a) and (c) hinting to a very weak shrinkage induced by HI in agreement with the results of the end-to-end distance. However, at the higher active force of $f^{a}=25$, MSIDs of wet active polymers are significantly lower than their dry counterparts, particularly for semiflexible polymers with $\kappa=10$, see Fig. 5 (c) and (d). The lower MSID curve of highly active wet polymers for $s>5$ indicates shrinkage at intermediate and large curvilinear distances. To further explore the geometrical features contributing to their more compact structure, we investigate their intrachain orientational correlation function.
@@
As can be seen from Fig. 7, for flexible chains with $\kappa=0$, the effective persistence lengths of wet and dry active chains are identical and almost independent of their activity levels. For semiflexible chains with $\kappa=10$, the effective persistence length $\ell_{p}$ of the wet chains are shorter than their dry counterparts. For dry active polymers, $\ell_{p}$ at the lowest activity ($f^{a}=0.01$) is close to the bare persistence length $\ell_{p}^{0}=10$ and that of the corresponding passive chains, see the red dashed line in Fig. 7. The effective persistence length of wet active chains at $f^{a}=0.01$ also agrees with the $\ell_{p}$ of passive wet polymers, as shown by the blue dashed line in Fig. 7. However, its value for the wet passive polymers is slightly smaller than $\ell_{p}^{0}$, reflecting the minor but visible effects of hydrodynamic interactions on the persistence length in the passive limit. The effective persistence lengths decrease with activity for both dry and wet chains, because high active forces increase the likelihood of the polymer backbone overcoming bending energy barriers to adopt more flexible conformations. Consistent with the results of the end-to-end distance, the $\ell_{p}$ of wet semiflexible chains exhibits a significant decrease at high activity levels. For instance, at a high active force of $f^{a}=25$, the effective persistence length of wet semiflexible chains matches that of flexible wet active polymers with $\kappa=0$. This finding, along with the end-to-end distance results, underscore the significant impact of HI on the conformation of highly active semiflexible chains.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a) $t=0$(b) $t=30$(c) $t=60$(d) $t=90$(e) $t=120$(f) $t=150$Refer to caption0.00.00.00.00.20.20.20.2Average fluid velocity
@@
Figure 8: Snapshots of wet active polymer with $\kappa=10$ and activity of $f^{a}=25$ at different times, showing the average flow field at monomer positions and their vicinity. The green segment shows the head of the polymer and the rest of the monomers are shown with yellow color. The color of the arrows encodes the strength of the flow field.
III.3 Investigation of mechanism of shrinkage
@@
As we have already seen from Figs. 2(g) and 3, HI lead to marked shrinkage of semiflexible polymers at high activity levels. To understand the underlying mechanism of shrinkage, we investigate the flow field in the vicinity of the backbone of an active chain with $\kappa=10$ and $f^{a}=25$, initialized as an straight line.
@@
Fig. 8 shows snapshots of this chain and the flow field at the monomer positions and their vicinity in a time series (see video S9). The flow field is measured as the average velocity of MPC particles within a sphere of radius $3$ centered at the position of interest. At $t=0$ when the chain is in a rod-like conformation, we observe a flow field nearly parallel to the backbone but heterogeneous in strength due to intrinsic fluid fluctuations. This heterogeneity in flow field strength, combined with entropic effects, renders the straight conformation unstable. In this example, the flow is stronger near the tail compared to the front segments. Because the back segments move faster than the front segments, they create a compressional flow which pushes the rear segemnts towards the front ones resulting in chain buckling, as illustrated in Fig.8(c) and (d). Finally, the difference in fluid velocity between the buckled and straight regions of the polymer further compresses its conformation until the entire structure becomes helix-like, as seen in Fig. 8(e) and (f). The general pattern in the motion of semiflexible wet active polymers consists of winding helix-like structures induced by flow heterogeneity, which creates local compressional flows and buckles the active polymer locally and create helix-like conformations. This is followed by unwinding of helix due fluctuations which make it unstable due to the resulting extensional flows, see video S8.
@@ IV.2 Translational dynamics
Figure 12: The longtime diffusion coefficient $D_{L}$ versus active force $f^{a}$ for chains with $\kappa=0$ and $10$. Close and open symbols represent data for chains with and without hydrodynamic interaction, respectively. The black dotted and dashed lines show the scaling slopes of $f^{a^{1}}$ and $f^{a^{0.5}}$. The dotted and dashed lines show the values at $f^{a}=0$ for $\kappa=0$ and $10$, respectively.
@@
Finally at larger lag times when $t\gg\tau_{e}$ the end-to-end vector (hence the total self propulsion) becomes uncorrelated to its initial direction and a final diffusion regime with enhanced diffusion coefficient emerges. The MSD curves at this regime follow the form $\langle\Delta\mathbf{\bm{R}}^{2}_{cm}(t\gg\tau_{e})\rangle=6D_{L}t$, where $D_{L}$ is the enhanced longtime diffusion coefficient. We extract the values of the diffusion coefficients by linear fits to the MSD curves at large lag times. The values of $D_{L}$ against activity are presented in Fig. 12 for wet and dry active polymers with bending stiffness values of $\kappa=0$ and $10$. In general, the wet chains exhibit more enhanced longtime dynamics. For dry active polymers, the $D_{L}$ increases linearly with activity in agreement with theoretical predictions Fazelzadeh et al. (2023); Philipps et al. (2022). Interestingly for flexible wet chains with $\kappa=0$, even with incorporation of HI, the $D_{L}$ increases linearly with activity. However for semiflexible wet active polymers, while at lower activity levels ($f^{a}<1$) the scaling behaviour of $D_{L}$ is linear in activity, at higher levels of activity where $f^{a}\geq 2.5$, another scaling regime emerges where $D_{L}\sim f^{a^{0.5}}$. The emergence of a new scaling regime at high active forces can be attributed to the HI-induced dramatic shrinkage of semiflexible wet chains, as visible from Fig. 3. Since the total active force is proportional to the end-to-end distance, the steep drop in $\sqrt{\langle R_{e}^{2}\rangle}$ decreases the scaling exponent of $D_{L}$ with respect to $f^{a}$.
@@ V Conclusions
In conclusion, our work offers novel insights into the interplay between activity and fluid-mediated interactions on the conformation and dynamics of tangentially-driven active polymers for which active force and conformation are coupled to each other. By implementing active chain hydrodynamics using the MPC fluid method, we pave the way for future investigations into the dynamics of tangentially-driven polymers in complex environments, such as confinement and crowded media.

### v3

@@ II.5 Simulation parameters
We choose the monomer diameter $\sigma$ as the unit of length, $k_{B}T$ as the unit of energy and $\tau=\sqrt{m\sigma^{2}/k_{B}T}$ as the unit of time. This leads to a dimensionless active force $f^{a}=F^{a}\sigma/k_{B}T$, which can also be interpreted as monomeric Peclet number Bianco et al. (2018). The wet polymers are simulated in a three-dimensional box with sides of length $L_{box}=60$ and periodic boundary conditions. The simulation box is split into collision cells of size $L_{cell}=1\sigma$. The mass of the MCPD fluid particles is set to $m=1$ and their collision time step is set to $\delta t=0.01\sqrt{m\sigma^{2}/k_{B}T}$. The stochastic rotation angle is chosen as $\alpha=130^{\circ}$. The number density of the MCPD particles is $\langle N_{c}\rangle=10$ leading to a total of $N_{s}=2.16\times 10^{6}$ MCPD particles in the simulation box.
The choice of $\delta t$, $N_{c}$ and $\alpha$ affect the viscosity of the resulting MPCD fluid $\eta=\eta_{k}+\eta_{c}$, which includes kinetic $\eta_{k}$ and collisional $\eta_{c}$ contributions Kikuchi et al. (2003); Winkler and Huang (2009). The kinetic viscosity $\eta_{k}$ is directly proportional to $\delta t$ Kikuchi et al. (2003) and is dominant at large timesteps representing gas-like behaviour of a fluid, whereas the collision viscosity $\eta_{c}$ is inversely proportional to $\delta t$. The latter predominates the total viscosity for small time steps mimicking viscous fluid behaviour Ripoll et al. (2005); Huang et al. (2015). It is given by Kikuchi et al. (2003)
$\eta_{c}=\frac{\langle N_{c}\rangle\;m}{18\;L_{cell}\;\delta t}\left(1-\cos(\alpha)\right)\left(1-\frac{1}{\langle N_{c}\rangle}\right)$ (18)
Our choice of MPCD parameters above yields $\eta_{c}\gg\eta_{k}$, hence the viscosity is $\eta\approx\eta_{c}\approx 82.1\frac{\sqrt{mk_{B}T}}{\sigma^{2}}$.
@@
We simulate 150 individual wet polymers. For each polymer, we first generate a three dimensional random walk of length $N_{m}$ and step size $\ell$. We use this sequence as the initial positions of the monomers. Once the polymer is created, the system is simulated for $10^{7}$-$10^{8}$ time steps to reach a steady state depending on the value of $f^{a}$. Then we run a production simulation of $2\times 10^{7}$ $2\times 10^{8}$ time steps, depending on the activity level. To ensure that our simulation results are not affected by the chosen box size and potential hydrodynamic self-interactions, we also ran additional simulations with bigger box size $L_{box}=100$ for $f^{a}=0.1$ and 25 for the semiflexible case, which led to identical static and dynamical features as discussed in more details in Appendix.B.
Simulations of dry active polymers, conducted for comparison with their wet counterparts, involve 120 individual dry active polymers described by Eq. (17) in a box with size $L_{box}=200$, while all other parameters being identical to those of wet active polymer model. The friction coefficient of monomers in dry active polymer simulation is chosen to be $\gamma=3\pi\eta\sigma\approx 773$. This leads to a ratio of $M/\gamma\approx 0.013$ which ensures that the dynamics of monomers are overdamped and inertial effects Fazelzadeh et al. (2023) remain negligible. To reach the steady state, we use $10^{7}$ to $10^{9}$ time steps, depending on the activity level. For the production run, twice the number of time steps used to reach the steady state are used. The time step for dry simulations is $\Delta t^{\prime}=0.002$.
@@ III Conformational features of wet active polymers
Figure 4: The probability distribution function of the end-to-end distance $P(R_{e})$ for different active forces $f^{a}$ for chains with $\kappa=0$ and $10$. Solid and dashed lines represent data for chains with and without hydrodynamic interaction, respectively. The insets show the standard deviations (STD) with respect to activity. The horizontal solid and dashed lines in the insets show the values at $f^{a}=0$ for wet and dry polymers respectively.
@@ III.1 End-to-end distance
To have a quantitative assessment of hydrodynamic effects on the mean size of active polymers, we examine the root of the mean squared end-to-end distances of the chains, i.e., $\sqrt{\langle R_{e}^{2}\rangle}=\sqrt{\langle|\mathbf{\bm{R}}_{N_{m}}-\mathbf{\bm{R}}_{1}|^{2}\rangle}$. Fig. 3 depicts the $\sqrt{\langle R_{e}^{2}\rangle}$ against activity for dry and wet chains with $\kappa=0$ and $10$, where the dotted and dashed lines show the corresponding $\sqrt{\langle R_{e}^{2}\rangle}$ of passive polymers. First, we note that for the passive and very low activity limit $f^{a}=0.01$, the $\sqrt{\langle R_{e}^{2}\rangle}$ of wet and dry polymer are equal within the error bars, confirming that for HI interactions do not affect static features of passive polymers. Upon increasing active force, for flexible chains with $\kappa=0$, we observe slightly smaller mean end-to-end distance values for wet polymers compared to their dry counterparts. As the activity level increases, the difference between the two becomes more pronounced. Looking into the $\sqrt{\langle R_{e}^{2}\rangle}$ of semiflexible chains with $\kappa=10$ in Fig. 3, we observe a more notable shrinkage of chains compared to their flexible counterparts, specially at larger active forces $f^{a}>1$ where the mean size of wet active polymers is about twice smaller than the dry ones.
To better assess the impact of HI on the conformation of active polymers, we examine the probability distribution function (PDF) of $R_{e}$, presented in Fig. 4 with the insets showing the corresponding standard deviation (STD) of each PDF. For both flexible and semiflexible polymers, the distribution of the end-to-end distance of wet active polymers shifts towards smaller $R_{e}$ values relative to their dry counterparts and the $P(R_{e})$ of wet active chains become narrower as their STD drops with increasing activity. This effect is particularly noticeable in the probability distribution $P(R_{e})$ of wet active polymers at $f^{a}=25$. In both flexible and semiflexible chains, narrow peaks emerge at lower $R_{e}$ values for wet active polymers, contrasting sharply with their dry counterparts. In the case of semiflexible chains, the HI-induced shrinkage of conformation is substantial enough to alter the overall shape of wet active polymers from a polymer-like conformation similar to that of dry polymers into a helix-like conformation, as demonstrated in Fig. 2.
@@ III.2 Interachain correlations
Having investigated the effects of HI on the statistical properties of end-to-end distance of active polymers, we study its impact on intrachain correlations by examining the bond-bond correlation function.
@@
As can be seen from Fig. 6, for fully flexible chains with $\kappa=0$, bond-bond correlations decay very rapidly as expected and the effective persistence lengths of wet and dry active chains are of the order of unity independent of their activity levels. For semiflexible chains with $\kappa=10$, the effective persistence lengths $\ell_{p}$ of the wet polymers become shorter than those of their dry counterparts as the activity increases. For both wet and dry chains, the $\ell_{p}$ values at $f^{a}=0$ and at the lowest activity ($f^{a}=0.01$) are close to the bare persistence length $\ell_{p}^{0}=10$, see the dashed lines in Fig. 6 for the $\ell_{p}$ of passive chains. The effective persistence lengths decrease with activity for both dry and wet chains, because high active forces increase the likelihood of the polymer backbone overcoming bending energy barriers to adopt more flexible conformations. Consistent with the results of the end-to-end distance, the $\ell_{p}$ of wet semiflexible chains exhibits a significant decrease at high activity levels. For instance, at a high active force of $f^{a}=25$, the effective persistence length of wet semiflexible chains matches that of flexible wet active polymers with $\kappa=0$. This finding, along with the end-to-end distance results, underscore the significant impact of HI on the conformation of highly active semiflexible chains.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a) $t=0$(b) $t=30$(c) $t=60$(d) $t=90$(e) $t=120$(f) $t=150$Refer to caption0.020.020.020.020.20.20.20.2Average fluid speed
@@
Figure 7: Snapshots of wet active polymer with $\kappa=10$ and activity of $f^{a}=25$ at different times, showing the average flow field at monomer positions and their vicinity. The green segment shows the head of the polymer and the rest of the monomers are shown in yellow. The color of the arrows encodes the speed of the flow field.
III.3 Enhancement of shrinkage with HI
@@
As we have already seen from Figs. 2(g) and 3, HI lead to greater shrinkage of semiflexible polymers at high activity levels compared to that observed in their dry counterparts. This feature is a manifestation of non-linear coupling of bending elasticity, activity and hydrodynamic effects for non-equilibrium polymers. To understand the underlying mechanism of the enhanced shrinkage, we investigate the flow field in the vicinity of the backbone of an active chain with $\kappa=10$ and $f^{a}=25$, initialized as an straight line, where the springs have their equilibrium rest length.
Fig. 7 shows snapshots of this chain and the smoothed coarse-grained flow field at the monomer positions and their vicinity in a time series (see video S9). The instantaneous smoothed flow field of a point of interest $P=(x,y,z)$ at any time is measured by averaging over the velocity of all fluid particles within a sphere of radius $3L_{cell}$ centered at $P$. At $t=0$ when the chain is in a rod-like conformation, since the activity is high we observe a strong flow field nearly parallel to the backbone but heterogeneous in strength with the given spatial distribution of active forces, see Eq. (10). The initial flow field for this conformation can be evaluated by summing the flow fields of stokeslets of equal strengths, apart from the end head and tail segments, which are all parallel to the chain direction, see Appendix. C. Doing so we find that the flow field is largest in the middle of the straight chain where active forces are distributed on both sides of the monomers. In contrast, at the head and tail segments, where forces are distributed only on one side of the monomer, the flow field is weakest.
@@
This heterogeneous distribution of flow field along the polymer backbone renders the straight conformation mechanically unstable. Consequently, the head segment moves slower than the rest of the polymer, while the tail does not experience the same slowdown, as it is pulled forward by the faster-moving segments ahead of it. This hydrodynamics instability coupled to dynamical active force rule, bending elasticity and thermal fluctuations lead to evolution of the conformation as presented in Fig. 7(b)-(f). Because the rear segments move faster than the front segments, they create a compressional flow which pushes the rear segments towards the front ones resulting in chain buckling, as illustrated in Fig. 7(c) and (d). Finally, the difference in fluid velocity between the buckled and straight regions of the polymer further compresses its conformation until the entire structure becomes coiled, as seen in Fig. 7(e) and (f). The general pattern in the motion of semiflexible wet active polymers consists of activity-induced flow heterogeneity generates compressional flows, causing the active polymer to buckle, leading to local folding and coiling. This is followed by flow fluctuations at the head segment, allowing the head to escape the buckled conformation, resulting in the gradual unwinding of the coiled structure, see video S8.
@@ IV.2 Translational dynamics
Figure 11: The longtime diffusion coefficient $D_{L}$ versus active force $f^{a}$ for chains with $\kappa=0$ and $10$. Close and open symbols represent data for chains with and without hydrodynamic interaction, respectively. The black dotted and dashed lines show the scaling slopes of $f^{a^{1}}$ and $\sqrt{f^{a}}$. The dotted and dashed lines show the values at $f^{a}=0$ for $\kappa=0$ and $10$, respectively.
@@
Finally at larger lag times when $t\gg\tau_{e}$ the end-to-end vector (hence the total self propulsion) becomes uncorrelated to its initial direction and a final diffusion regime with enhanced diffusion coefficient emerges. The MSD curves at this regime follow the form $\langle\Delta\mathbf{\bm{R}}^{2}_{cm}(t\gg\tau_{e})\rangle=6D_{L}t$, where $D_{L}$ is the enhanced longtime diffusion coefficient. We extract the values of the diffusion coefficients by linear fits to the MSD curves at large lag times. The values of $D_{L}$ against activity are presented in Fig. 11 for wet and dry active polymers with bending stiffness values of $\kappa=0$ and $10$. In general, the wet chains exhibit more enhanced longtime dynamics. For dry active polymers, the $D_{L}$ increases linearly with activity in agreement with theoretical predictions Fazelzadeh et al. (2023); Philipps et al. (2022). Interestingly for flexible wet chains with $\kappa=0$, even with incorporation of HI, the $D_{L}$ increases linearly with activity. However for semiflexible wet active polymers, while at lower activity levels ($f^{a}<1$) the scaling behaviour of $D_{L}$ is linear in activity, at higher levels of activity where $f^{a}\geq 2.5$, another scaling regime emerges where $D_{L}\sim\sqrt{f^{a}}$. The emergence of a new scaling regime at high active forces can be attributed to the HI-induced dramatic shrinkage of semiflexible wet chains, as visible from Fig. 3. Since the total active force is proportional to the end-to-end distance, the steep drop in $\sqrt{\langle R_{e}^{2}\rangle}$ decreases the scaling exponent of $D_{L}$ with respect to $f^{a}$.
@@ V Conclusions
In conclusion, our work offers novel insights into the interplay between activity and fluid-mediated interactions on the conformation and dynamics of tangentially-driven active polymers for which active force and conformation are coupled to each other. By implementing active chain hydrodynamics using the MCPD fluid method, we pave the way for future investigations into the dynamics of tangentially-driven polymers in complex environments, such as confinement and crowded media.